OpenStudy (kimberly_pr):

Solve for x (5/6)x = 30

Arcadiouse (arcadiouse):

x = 24.9?

OpenStudy (will.h):

if it is like the following \[\frac{ 5 }{ 6 }x = 30\] then you can just divide both sides on 5/6

Arcadiouse (arcadiouse):

5/6 = .83 30 x .83 = 24.9

OpenStudy (kimberly_pr):

Thanks @Arcadiouse :) but I think they want it in fraction form. And ohhh I see @Will.H thank you so mcuh

Arcadiouse (arcadiouse):

24 9/10

OpenStudy (kimberly_pr):

SO it would be 30/1 / 5/6 and I have to find their least common denominator before dividing right? @Will.H

OpenStudy (kimberly_pr):

SO the answer would be x=6 right? That's what I got @Will.H

OpenStudy (will.h):

if you divide 30 on 5/6 the answer youd get is 36 and to check that if correct or not you should substitute x = 36 in the original equation so therefor \[\frac{ 5 }{ 6 } * 36 = 30\] \[30 = 30\] \ which confirms that x = 36 is a correct answer for the equation hope that helps

OpenStudy (kimberly_pr):

oh ok, thank you can i ask one more question? @Will.H

OpenStudy (will.h):

in a separate post. You're welcome